# The temporary impact of permanent hiring incentives: Evidence from Italy

### v1

@@ 1 Introduction
To mitigate these potential negative consequences without reversing the liberalization of temporary contracts, policymakers of advanced economies, like Italy, Spain, and France, [^1] Among others: for Spain law 43/2006, Tarifa Joven (2014-2016); for France Contrat nouvelle embauche (CNE) – 2005, Régime des emplois francs – da gennaio 2020. See Section 2 for further details on the Italian incentives. have frequently implemented incentives to encourage firms to hire workers under open-ended contracts. These policies generally entail bonuses, most commonly payroll tax reductions or cuts in social security contributions, granted to firms that either hire new employees under permanent contracts or convert temporary contracts into permanent ones.
@@
Our study focuses on the social contribution exemption introduced in Italy over the course of the mid-2010s to employers hiring young workers into new open-ended contracts (OECs) or converting existing fixed-term contracts to permanent positions. Over the years, the scheme has been subject to several changes in its eligibility criteria. [^2] A detailed discussion is presented in Section 2. In particular, we take advantage of the unique exogenous discontinuity in treatment eligibility that was introduced by the 2018 Budget Law to identify the causal impact of incentives on contract conversions by using administrative worker-level data for Tuscany. With respect to the existing literature, we concentrate on incentives specific to permanent contracts, studying their effects over a target population of workers not detached from the labor market, and avoid confounding these effects with those of other concurrent incentives for temporary hires. We provide a comprehensive assessment of the reform across four key dimensions: the immediate effect on contract conversions, the medium-term impact on beneficiaries’ career trajectories, the heterogeneous effects of the policy, and the substitution patterns between eligible and ineligible cohorts.
@@
The literature on the short-term effects of hiring subsidies often documents moderately positive results on employment among eligible workers during the subsidy period. For example, a generous one-shot wage subsidy in Belgium for unemployed youth (“Win-Win Plan” after the Great Recession) raised the likelihood of finding private-sector employment by about 10 percentage points in the first year for eligible young people (Albanese et al.,, 2024). In France, a one-year emergency hiring credit targeting low-wage jobs during the Great Recession led to significant employment gains (Cahuc et al.,, 2019), with no measurable increase in wages. Similarly, a large payroll tax cut for young workers in Sweden (2007–2009) led to a 2–3 percentage point increase in youth employment for the targeted age group, with treated firms hiring more young workers and expanding their overall workforce (Saez et al.,, 2019). All these incentives targeted first-time or disadvantaged low-pay workers, encompassing both permanent and temporary contracts. These groups of beneficiaries may differ significantly from the general working population, limiting the external validity of their results. By contrast, the hiring incentives introduced in Italy over the past decade have been designed primarily to promote more stable forms of employment within the general workforce. The literature, so far, has focused on the evaluation of the 2012 (Ciani and De Blasio,, 2015) and 2015 (see, among others, Sestito and Viviano,, 2018; Boeri and Garibaldi,, 2019; Deidda et al.,, 2021; Brunetti et al.,, 2022; Ardito et al.,, 2023; Santoni et al.,, 2024) schemes, generally reporting a positive impact. However, these incentives often overlapped with broader schemes for younger workers that also covered temporary contracts, making it difficult to isolate the specific contribution of permanent hiring subsidies. The 2018 reform, by introducing an age threshold that extended eligibility to a group otherwise excluded from any other incentives, provides a unique opportunity to disentangle this effect.
@@
In addition to these immediate impacts, some of the literature has also focused on the persistence of the employment gains from hiring incentives long after their ending, that is, the medium-term effects in terms of job stability and career prospects. The evidence here is mixed and often less encouraging. In the context of the aforementioned Belgian youth subsidy, the positive impact for the least skilled youth was strictly short-lived: high school dropouts did not see lasting improvement beyond the subsidy period (Albanese et al.,, 2024). In contrast, high school graduates in the same program experienced some persistent benefits, even if modest in scale, in terms of job-finding probability and employment spells. Similarly, Saez et al., (2021) studied the longer-term impacts of the 2007 incentives in Sweden, finding significantly positive employment effects for the subsequent career of the beneficiaries, a result that contrasts with the findings of (Egebark and Kaunitz,, 2018), who found little evidence of lasting effects for the treated individuals once they aged out of eligibility. Moreover, (Sjögren and Vikström,, 2015), analyzing another Swedish policy, and (Batut,, 2021), examining a hiring credit for small firms in France, found that although employment declines once subsidies expire, workers who obtained subsidized jobs still exhibit a higher probability of being employed even after the expiration of the subsidies. Differently, (Desiere and Cockx,, 2022) studied the abolition of a hiring subsidy targeted at older long-term unemployed jobseekers in Belgium and found that incentives mainly create temporary, short-lived employment. Narrowing the focus to open-ended-contract incentives, Ardito et al., (2025) analyze the 2015 Italian incentive scheme, which granted a full rebate of social security contributions for three years to firms hiring workers on new permanent contracts. Their results show that, although the policy initially reduced separation risks, this protective effect was short-lived, with a sharp spike in exits occurring precisely when the subsidy expired, and no evidence of any lasting improvement in either job or employment security. As with the short-term evidence, however, the longer-term effects on open-ended contracts remain difficult to disentangle because of overlapping schemes or their focus on specific populations, or both.
@@
Lastly, another very relevant aspect to consider when evaluating employment incentives is the possible spillover effects and substitution that can undermine the impact. A key concern is that subsidizing some jobs or workers may simply displace employment opportunities elsewhere. The available evidence, which mostly concerns first-time labor market entrants, generally points to negative or negligible spillovers onto non-eligible workers. For example, the French low-wage hiring credit of 2009 showed no displacement of incumbent workers or non-eligible hires at treated firms (Cahuc et al.,, 2019). Likewise, evaluations of youth-focused subsidies in Belgium detect minimal substitution away from slightly older (ineligible) workers (Albanese et al.,, 2024). Furthermore, Ciani and De Blasio, (2015), analyzing the 2012 Italian incentives, found no substitution effects. The same limitations noted earlier also restrict the generalizability of these studies to the effects on open-ended contracts.
@@
In short, a comprehensive evaluation of hiring incentives for the creation of new open-ended contracts, one that jointly considers short- and medium-term effects as well as possible substitution effects, remains missing from the literature. We identify the policy effects by using a combination of difference-in-differences and regression discontinuity designs. We find that 2018 incentives effectively increased the immediate conversion rate from temporary to open-ended contracts for eligible workers. To account for overall substitution effects with slightly older peers, we perform placebo tests in the treatment year, obtaining null results. We also find no within-firm substitution effects as the change in the proportion of eligible workers around the cutoff with respect to the previous year does not appear to significantly affect the conversion rate of eligible workers. However, we do not detect any significant longer-term effect one, two, three, and four years after the introduction of the policy; indeed, the incentives’ effects disappear within a year of the reform. A possible explanation is that the policy merely anticipated the conversion to a permanent contract for temporary workers who would have been converted anyway.
@@ 2 Institutional Background
The first set of these incentives was introduced as part of the 2012 Monti-Fornero reform (Law 92/2012), among other labor market reforms meant to regulate temporary contracts and reduce dismissal costs under specific circumstances. The incentives initially encompassed the hiring, under both permanent and temporary contracts, of people over 50 years old and disadvantaged women. [^3] Were included: women of any age, residing anywhere, who have been without regular paid employment for at least 24 months or if they were living in disadvantaged areas or working in sectors with high gender employment disparities, the limit was lowered to 6 months. The incentives scheme entitled the employer to a 50% reduction in the social contribution payment for the hiring of eligible workers for up to 18 months, provided that it results in a net increase in the company’s workforce. These incentives, which still remain in effect, were accompanied by an extraordinary incentive scheme [^4] Decree October 5th 2012 which included men below 30 years old and women of all ages, and granted employers an exemption ranging from EUR 3,000 for short temporary contracts to EUR 12,000 for conversion toward permanent contracts. However, given the limited amount of available funds, this subsidy was depleted in just a few days. Ciani and De Blasio, (2015) examined the immediate effects of the extraordinary scheme by focusing on its effect on temporary-to-permanent transformations.
@@
The biennium 2014–2015 marked a period of significant reforms in Italy, many of which were included in the package known as the Jobs Act. The major changes sought to incentivize hiring by making the Italian labor market more flexible, and entailed the reduction of firing costs for workers with OECs, which was complemented by the introduction of generous hiring incentives. This incentive scheme consisted of a three-year exemption from the full amount of social security contributions for all open-ended contract hires and transformation from fixed-term ones made during that year, with no age restrictions, for a maximum of EUR 8,000. The entanglement between all these measures, including the incentives included in the Youth Guarantee, has complicated the evaluation of the Jobs Act. Nevertheless, the studies focusing on such evaluation (Sestito and Viviano,, 2018; Boeri and Garibaldi,, 2019; Deidda et al.,, 2021; Brunetti et al.,, 2022; Ardito et al.,, 2023; Santoni et al.,, 2024) found small to moderate effects, mostly attributable to the incentives rather than to the reduction of firing costs. Initially, this incentive was intended to cover both 2015 and 2016, but with the 2016 Budget Law (L. 208/2015), the exemption from social security contributions payment was changed to a 40% reduction for the hirings and transformations in 2016, for a maximum duration of 2 years.
@@
Finally, the 2018 budget law (Law n. 205/2017) updated the employment incentives for new and converted permanent hires, keeping the incentives for temporary contracts unchanged. While the generosity of the former incentives was reduced, entailing a 50% reduction in social contributions (up to a maximum annual amount of EUR 3,000), the duration was increased to 36 months, and the age limit threshold was raised to 35 years old. This change in cohort eligibility was intended to be a temporary one, and was meant to return to 30 for the subsequent years, 2019 and 2020. However, at the end of 2019, with the new Budget Law for 2020 (Law n.160/2019), the threshold was retroactively set back to 35 for 2019 and 2020. While, technically, employers were not aware of the retroactive eligibility change until the end of that year, it is unclear if they were actually able to anticipate it.
2018 was also marked by a significant labor market reform. In July, the so-called Decreto Dignità was approved, marking a partial reversal of the Jobs Act and responding to a Constitutional Court ruling that had struck down parts of its dismissal rules. The reform reduced the maximum duration of temporary contracts from 36 to 24 months, lowered the maximum number of extensions from five to four, and reinstated the requirement to provide a justification for contracts exceeding 12 months. Although this policy did not directly affect the incentives under study, tighter regulations on temporary contracts may have increased the likelihood of permanent conversions, especially for workers already close to the new maximum duration. Since the reform became effective only in November 2018, almost the entire year remained unaffected, allowing us to estimate the immediate effect without major drawbacks. Moreover, the 35-year age threshold was not affected by the reform, meaning that workers around the cutoff were treated uniformly. Our medium-run evaluation is therefore not biased by this institutional change.
An additional discussion concerns the implementation of these schemes and their restrictions. With the exception of the 2012 extraordinary incentives, the implementation of the incentives has been almost automatic for employers. In practice, firms applied the exemptions directly when calculating their social security contributions and reported them through the standard INPS system, later consolidated in their annual tax filings. An additional restriction has been introduced since 2017: employers could not benefit from hiring incentives if they had carried out dismissals in the same production unit during the preceding six months. This “anti-layoff” clause was meant to prevent abuses in which subsidized youth hires were immediately offset by terminations of existing staff.
@@ 3 Data
These communications fall into four categories: hiring, extensions (proroghe), transformations (trasformazioni) [^7] I.e., tenure changes from temporary to permanent., and terminations (cessazioni). A detailed set of worker and job-specific characteristics is also included with each communication. Together, these mandatory communications allow us to reconstruct labor flows that have occurred since the system was implemented, covering around 12 million job spells in Tuscany alone. The dataset links job contracts with unique worker and firm identifiers, allowing us to track the evolution and, possibly, conversion of temporary contracts over time.
To best study conversions, we start by considering the set of mandatory communications concerning standard temporary contracts [^8] Standard temporary contracts, vis-a-vis other atypical ones, represent by a wide margin the most common form of fixed-term contract in the Italian labour market and are the only ones that can be converted to open-ended contracts. Accordingly, atypical contracts are not considered in our analysis. for each year of our period of interest, ranging from 2014 to 2019. For each year, we select all the job relationships of people with a standard temporary contract who are aged 34 to 36 years old and who have never held a permanent contract before.
@@ 3.1 Descriptive evidence on policy take-up for conversions
Some stylized facts immediately emerge from the figure. Firstly, both the 2015 and 2018–2019 windows are characterised by significant overall increases in conversions. Nonetheless, the trends in conversions appear to be parallel among the age groups. A single discontinuity emerges among the cohorts, appearing only in 2018, further motivating our analysis. In that year, conversions rose across all age groups, but the increase was noticeably smaller for those aged over 35. Interestingly, the reintroduction of hiring incentives in 2017 was not accompanied by any differential increase in conversions for the eligible population of under-30s. Possibly, the combination of incentives for temporary and permanent hires for the same age group was detrimental to permanent conversions.
in
Figure 2: Transformations by age-groups
@@
Figure 3 further disentangles these trends by looking at individuals turning 34, 35, and 36 in each year. Even after these sample restrictions, the gap between the 34- and 35-year-old groups increased significantly in 2018. This gap narrowed significantly in 2019, consistent with the retroactive increase in the age limit from 30.
Figure 3: Transformations by age close to the policy threshold

### v2

# The temporary impact of permanent employment incentives: Evidence from Italy

@@ 1 Introduction
To mitigate these potential negative consequences without reversing the liberalization of temporary contracts, policymakers of advanced economies, like Italy, Spain, and France, [^1] Among others: for Spain law 43/2006, Tarifa Joven (2014-2016); for France Contrat nouvelle embauche (CNE) – 2005, Régime des emplois francs – 2020. See Section 2 for further details on the Italian incentives. have frequently implemented incentives to encourage firms to increase job stability. These policies typically involve bonuses, often in the form of payroll tax reductions or cuts in social security contributions, awarded to firms that either hire new employees under permanent contracts or convert temporary contracts into permanent ones.
@@
Our study focuses on the social contribution exemption introduced in Italy during the mid-2010s for employers hiring young workers into new open-ended contracts (OECs) or converting existing fixed-term contracts to permanent positions. Over the years, the scheme has undergone several changes to its eligibility criteria. [^2] A detailed discussion is presented in Section 2. In particular, we leverage the unique exogenous discontinuity in treatment eligibility introduced by the 2018 Budget Law to estimate the causal impact of incentives on contract conversions, utilizing administrative worker-level data for Tuscany. With respect to the existing literature, we concentrate on permanent contract conversions rather than new hires, allowing us to study the effects of these incentives over a target population of workers not detached from the labor market, and to avoid confounding these effects with those of other concurrent incentives for temporary hires. We provide a comprehensive assessment of the reform across four key dimensions: the immediate effect on contract conversions, the medium-term impact on beneficiaries’ career trajectories, the heterogeneous effects of the policy, and the substitution patterns between eligible and ineligible cohorts.
@@
The literature on the short-term effects of hiring subsidies often documents moderately positive results on employment among eligible workers during the subsidy period. For example, a generous one-shot wage subsidy in Belgium for unemployed youth (“Win-Win Plan” after the Great Recession) raised the likelihood of finding private-sector employment by about 10 percentage points in the first year for eligible young people (Albanese et al.,, 2024). In France, a one-year emergency hiring credit targeting low-wage jobs during the Great Recession led to significant employment gains (Cahuc et al.,, 2019), with no measurable increase in wages. Similarly, a large payroll tax cut for young workers in Sweden (2007–2009) led to a 2–3 percentage point increase in youth employment for the targeted age group, with treated firms hiring more young workers and expanding their overall workforce (Saez et al.,, 2019). All these incentives targeted first-time or disadvantaged low-paid workers, encompassing both permanent and temporary contracts. These groups of beneficiaries may differ significantly from the general working population, limiting the relevance of their results in contexts where workers are already in temporary positions.
By contrast, the hiring incentives introduced in Italy over the past decade have been designed primarily to promote more stable forms of employment within the general workforce. The literature, so far, has focused on the evaluation of the 2012 (Ciani and De Blasio,, 2015) and 2015 (see, among others, Sestito and Viviano,, 2018; Boeri and Garibaldi,, 2019; Deidda et al.,, 2021; Brunetti et al.,, 2022; Ardito et al.,, 2023; Santoni et al.,, 2025) schemes, generally reporting a positive impact. However, these incentives often overlapped with broader schemes for younger workers that also covered temporary contracts, making it difficult to isolate the specific contribution of permanent hiring subsidies. The 2018 reform, by introducing an age threshold that extended eligibility to a group otherwise excluded from any other incentives, provides a unique opportunity to disentangle this effect.
@@
In addition to these immediate impacts, some literature has also focused on the longer-term persistence of employment gains from hiring incentives, often in terms of job stability and career prospects. This evidence is mixed and not very encouraging. In the context of the aforementioned Belgian youth subsidy, Albanese et al., (2024) found that the positive impact for high school dropouts did not persist beyond the subsidy period, while high school graduates in the same program experienced some modest yet persistent career benefits. Focusing instead on the abolition of another Belgian hiring subsidy targeted at older long-term unemployed, Desiere and Cockx, (2022) also found that incentives mainly create temporary, short-lived employment. In contrast, Saez et al., (2021) studied the longer-term impacts of the 2007 incentives in Sweden, finding significantly positive employment effects for the subsequent career of the beneficiaries, a result that contrasts with the findings of Egebark and Kaunitz, (2018), who found little evidence of lasting effects for the treated individuals once they aged out of eligibility. Moreover, Sjögren and Vikström, (2015), analyzing another Swedish policy, and Batut, (2021), examining a hiring credit for small firms in France, found that although employment declines once subsidies expire, workers who obtained subsidized jobs still exhibit a higher probability of being employed even after the expiration of the subsidies. Narrowing the focus to the 2015 Italian incentives, Ardito et al., (2025) found that although the policy initially reduced separation risks, this effect was short-lived, with a sharp spike in exits occurring precisely when the subsidy expired. As with the short-term evidence, these contrasting results may arise from the presence of overlapping schemes, the focus of incentives on specific populations, or both.
@@
Other strands of literature have focused on spillover and substitution effects. A key concern is that subsidizing some jobs or workers may simply displace employment opportunities elsewhere. The available evidence, primarily concerning first-time labor market entrants, generally indicates negative or negligible spillovers on non-eligible workers. For example, the French low-wage hiring credit of 2009 showed no displacement of incumbent workers or non-eligible hires at treated firms (Cahuc et al.,, 2019). Likewise, evaluations of youth-focused subsidies in Belgium detected minimal substitution away from slightly older (ineligible) workers (Albanese et al.,, 2024). Similarly, Ciani and De Blasio, (2015), analyzing the 2012 Italian incentives, found no substitution effects. The same limitations noted earlier also restrict the generalizability of these studies to the effects on open-ended contract conversions.
@@
In short, a comprehensive evaluation of hiring incentives that focuses on the transition from temporary to permanent contracts, and that disentangles these incentives from more general (applicable to any contract) and targeted (applicable to specific vulnerable groups) employment incentive schemes, remains missing from the literature. In our study, we identify the policy effects of the 2018 Italian incentives on contract transitions by using a combination of difference-in-differences and regression discontinuity designs. We find that the incentives effectively increased the immediate conversion rate from temporary to open-ended contracts for eligible workers. The effect is more pronounced for men, and for those working in the industrial and service sectors, in medium-skilled occupations, and in small firms. To account for overall substitution effects with slightly older peers, we perform placebo tests in the treatment year, obtaining null results. We also find no within-firm substitution effects as the change in the proportion of eligible workers around the cutoff with respect to the previous year does not appear to significantly affect the conversion rate of eligible workers. However, we do not detect any significant longer-term effect one, two, three, and four years after the introduction of the policy; indeed, the incentives’ effects, and their heterogeneities, disappear within a year of the reform. This evidence suggests that the policy merely anticipated the conversion to a permanent contract for temporary workers who would have been converted anyway.
@@ 2 Institutional Background
The first set of these incentives was introduced as part of the 2012 Monti-Fornero reform (Law 92/2012), among other labor market reforms meant to regulate temporary contracts and reduce dismissal costs under specific circumstances. The incentives initially encompassed the hiring, under both permanent and temporary contracts, of people over 50 years old and disadvantaged women. [^3] Were included: women of any age, residing anywhere, who have been without regular paid employment for at least 24 months or if they were living in disadvantaged areas or working in sectors with high gender employment disparities, the limit was lowered to 6 months. The incentives scheme entitled the employer to a 50% reduction in the social contribution payment for the hiring of eligible workers for up to 18 months, provided that it results in a net increase in the company’s workforce. These incentives, which still remain in effect, were accompanied by an extraordinary incentive scheme [^4] Decree October 5th 2012 which included men below 30 years old and women of all ages, and granted employers an exemption ranging from EUR 3,000 for short temporary contracts to EUR 12,000 for conversion toward permanent contracts. However, given the limited amount of available funds, this subsidy was depleted in just a few days. Ciani and De Blasio, (2015) examined the immediate effects of the extraordinary scheme by focusing on its effect on temporary-to-permanent transitions.
@@
The biennium 2014–2015 marked a period of significant reforms in Italy, many of which were included in the package known as the Jobs Act. The major changes sought to incentivize hiring by making the Italian labor market more flexible, and entailed the reduction of firing costs for workers with OECs, which was complemented by the introduction of generous hiring incentives. This incentive scheme consisted of a three-year exemption from the full amount of social security contributions for all open-ended contract hires and transition from fixed-term ones made during that year, with no age restrictions, for a maximum of EUR 8,000. The entanglement between all these measures, including the incentives included in the Youth Guarantee, has complicated the evaluation of the Jobs Act. Nevertheless, the studies focusing on such evaluation (Sestito and Viviano,, 2018; Boeri and Garibaldi,, 2019; Deidda et al.,, 2021; Brunetti et al.,, 2022; Ardito et al.,, 2023; Santoni et al.,, 2025) found small to moderate effects, mostly attributable to the incentives rather than to the reduction of firing costs. Initially, this incentive was intended to cover both 2015 and 2016, but with the 2016 Budget Law (L. 208/2015), the exemption from social security contributions payment was changed to a 40% reduction for the hirings and transitions in 2016, for a maximum duration of 2 years.
@@
Finally, the 2018 budget law (Law n. 205/2017) updated the employment incentives for new and converted permanent hires, keeping the incentives for temporary contracts unchanged. While the generosity of the former incentives was reduced, entailing a 50% reduction in social contributions (up to a maximum annual amount of EUR 3,000), the duration was increased to 36 months, and the age limit threshold was raised to 35 years old. This change in cohort eligibility was intended to be a temporary one, and was meant to return to 30 for the subsequent years, 2019 and 2020. However, at the end of 2019, with the new Budget Law for 2020 (Law n.160/2019), the threshold was retroactively set back to 35 for 2019 and then maintained for 2020 as well. While technically, employers were not aware of the retroactive eligibility change until the end of that year, it is unclear if they were actually able to anticipate it.
2018 was also marked by a significant labor market reform. In July, the so-called Decreto Dignità was approved, marking a partial reversal of the Jobs Act and responding to a Constitutional Court ruling that had struck down parts of its dismissal rules. Although this policy did not directly affect the incentives under study, leaving the cutoff and the generosity of the incentives unchanged, tighter regulations on temporary contracts may have increased the likelihood of permanent conversions, especially for workers already close to the new maximum duration. However, the reform only became completely effective in November 2018, allowing us to estimate the immediate effect without major drawbacks. Furthermore, since the 35-year age threshold was not affected by the reform, workers around the cutoff were treated uniformly, thereby maintaining the internal validity of our results. This being considered, in Appendix A, we show how the estimated short-term policy effect is unaffected by the introduction of this concurrent policy.
An additional discussion concerns the implementation of these schemes and their restrictions. With the exception of the 2012 extraordinary incentives, the implementation of the incentives has been almost automatic for employers. In practice, firms applied the exemptions directly when calculating their social security contributions and reported them through the standard INPS system, later consolidated in their annual tax filings. An additional restriction was introduced in 2017: employers could not benefit from hiring incentives if they had carried out dismissals in the same production unit within the preceding six months. This “anti-layoff” clause was meant to prevent abuses in which subsidized youth hires were immediately offset by terminations of existing staff.
@@ 3 Data
These communications fall into four categories: hiring, extensions (proroghe), conversions (trasformazioni), [^7] I.e, tenure changes from temporary to permanent. and terminations (cessazioni). A detailed set of worker and job-specific characteristics is also included with each communication. Together, these mandatory communications allow us to reconstruct labor flows that have occurred since the system was implemented, covering around 12 million job spells in Tuscany alone. The dataset links job contracts with unique worker and firm identifiers, allowing us to track the evolution and, possibly, conversion of temporary contracts over time.
To best study conversions, we start by considering the set of mandatory communications concerning standard temporary contracts [^8] Standard temporary contracts, vis-a-vis other atypical ones, represent the most common form of fixed-term contract in the Italian labour market. In fact, according to a report from the National Social Security Institution (INPS), standard temporary contracts represented 45% of the total number of new hires in 2018, while atypical contracts accounted for only 33%. Accordingly, atypical contracts are not considered in our analysis. Source: https://www.inps.it/content/dam/inps-site/pdf/dati-analisi-bilanci/osservatori-statistici/osservatorio-precariato/5328KEY-osservatorio_precariato_gen_dic_2018.pdf?utm_source=chatgpt.com for each year of our period of interest, ranging from 2014 to 2019. For each calendar year, we select all the job relationships on standard fixed-term contracts held by individuals aged 34–36 who have never previously held a permanent contract. In the next section, we look at a wider group of individuals in order to present some preliminary evidence.
@@ 3.1 Descriptive evidence on policy take-up for conversions
Several stylized facts are immediately evident from the figure. Firstly, both the 2015 and 2018–2019 windows are characterised by significant overall increases in conversions. Nonetheless, the trends in conversions appear to be parallel among the age groups. A single discontinuity emerges among the cohorts, appearing only in 2018, further motivating our analysis. In that year, conversions rose across all age groups, but the increase was noticeably smaller for those aged 35 and older. Interestingly, the reintroduction of hiring incentives in 2017 was not accompanied by any differential increase in conversions for the eligible population of under-30s. Possibly, the combination of incentives for temporary and permanent hires for the same age group was detrimental to permanent conversions.
Figure 2: Conversions by age-groups
@@
Figure 3 further disentangles these trends by looking at individuals turning 34, 35, and 36 in each year. Even after these sample restrictions, the gap between the 34- and 35-year-old groups increased significantly in 2018. The conversion for people aged 36 years old in 2018 increased, but at a lower pace. The previously mentioned gap narrowed significantly in 2019, which aligns with the fact that these individuals were initially excluded from coverage that year..
Figure 3: Conversions by age close to the policy threshold